# https://people.engr.tamu.edu/nimak/Papers/CoherentGS

### v1

@@ 2.2 Sparse Novel View Synthesis
Concurrent to our technique, Zhu et al. [62] (FSGS) and Xiong et al. [54] (SparseGS) propose to utilize 3DGS for sparse view synthesis. However, unlike our method, they do not present a way to enforce coherency between nearby Gaussians, producing occasional floaters. Additionally, these methods fill in the regions that are occluded in all the input images with elongated Gaussians producing blurry results. In contrast, our approach allows us to identify and inpaint the occluded regions and produce high-quality hallucinated details (see Fig. 1 - right).
@@ 4.1 Coherent 3D Gaussian Optimization
Single-view Constraint We enforce per-view smoothness by utilizing an implicit convolutional decoder. Specifically, as shown in Fig. 3, this decoder takes the view index $n$ as the input and estimates the residual depth for the entire image, i.e., $\Delta D_{n}=f_{\phi}(n)$, where $\phi$ refers to the parameters of the decoder. Instead of updating the residual depth at individual pixels, we perform the optimization by modifying the decoder parameters $\phi$. This ensures that the residual depth is smooth and object surfaces are coherently deformed during optimization.
However, smooth deformation means that the decoder struggles to handle the sharp depth discontinuities between the objects. We address this by obtaining a $C$-channel binary segmentation mask through dividing each image into $C$ (5 in our implementation) separate regions, each with similar depths. Specifically, we use the approach by Yang et al. [56] to estimate the monocular depth and follow the strategy by Wang et al. [50] to divide the image into $C$ regions based on the input depth (see Fig. 3 - right). The decoder in this case, also produces a $C$-channel residual depth. The final residual is obtained as $\Delta D_{n}=\text{chsum}(S\odot f_{\phi}(n))$, where chsum refers to the channelwise summation operation.
@@ 4.2 Additional Regularization
While 3DGS optimizes the objective by only sampling the center of each pixel, we found this strategy to be problematic for sparse input setting. In this case, the Gaussians will be deformed to match the color at the center of each pixel, leaving the remaining areas uncovered. As a result, the surfaces, when viewed from a novel view, will appear semi-transparent. To address this issue, we perform the optimization at multiple samples within each pixel. The multisampling ensures that the Gaussians properly cover each pixel, resulting in significantly improved images, as shown in the supplementary video and Table 3.
@@
Figure 6: Gaussian Scaling
@@
where $r$ is the radius of the sphere and is used to create the isometric scale matrix, and $f$ represents the vertical focal length of the input image with height $H$. For the color, represented in terms of the SH coefficients, we set the DC value to the pixel color and other coefficients to zero. Finally, we initialize the global opacity to a constant.
@@ 5.1 Comparisons
As discussed previously, unlike other methods that fill in the occluded regions with blurry and repetitive textures, our method does not reconstruct these areas. This gives us a unique advantage as we can identify and inpaint these areas to hallucinate high-quality details. As a proof of concept, we use a simple strategy to inpaint these areas with a diffusion model and project them to 3D using monocular depth. As shown in Fig. 1 (right) the hallucinated details are of high-quality and consistent (see supplementary video).
@@ 7 Implicit Decoder Architecture
We use the convolutional decoder network as proposed by Bemana et al. [5, 36]. The input coordinate is concatenated to the coordconv [30] layer in the first layer of the decoder. This is followed by a series of convolutional and upsampling layers (bi-linear interpolation) until the desired output resolution is reached. The number of parameters is controlled using a scalar capacity factor that is multiplied to a preset of each layer, e.g., [2c, 4c, 8c] where c is the capacity factor. We use capacity factor [10, 15, 18] for the depth decoder corresponding to 2, 3 and 4 inputs, respectively. Similarly, we use [6, 10, 12] for the opacity decoder.

### v2

@@ 2.2 Sparse Novel View Synthesis
Concurrent to our technique, Zhu et al. [68] (FSGS), Xiong et al. [57] (SparseGS) and Li et al. [27] (DNGaussian) propose to utilize 3DGS for sparse view synthesis. However, unlike our method, they do not present a way to enforce coherency between nearby Gaussians, producing occasional floaters. Additionally, these methods fill in the regions that are occluded in all the input images with elongated Gaussians producing blurry results. In contrast, our approach allows us to identify and inpaint the occluded regions and produce high-quality hallucinated details (see Fig. 1 - right).
@@ 4.1 Coherent 3D Gaussian Optimization
Single-view Constraint We enforce per-view smoothness by utilizing an implicit convolutional decoder. Specifically, as shown in Fig. 3, this decoder takes the normalized view index $n$ as the input and estimates the residual depth for the entire image, i.e., $\Delta D_{n}=f_{\phi}(n)$, where $\phi$ refers to the parameters of the decoder. Instead of updating the residual depth at individual pixels, we perform the optimization by modifying the decoder parameters $\phi$. This ensures that the residual depth is smooth and object surfaces are coherently deformed during optimization.
However, smooth deformation means that the decoder struggles to handle the sharp depth discontinuities between the objects. We address this by obtaining a $C$-channel binary segmentation mask, $S$, through dividing each image into $C$ (5 in our implementation) separate regions, each with similar depths. Specifically, we use the approach by Yang et al. [60] to estimate the monocular depth and follow the strategy by Wang et al. [53] to divide the image into $C$ regions based on the input depth (see Fig. 3 - right). The decoder in this case, also produces a $C$-channel residual depth. The final residual is obtained as $\Delta D_{n}=\text{chsum}(S\odot f_{\phi}(n))$, where chsum refers to the channelwise summation operation.
@@ 4.2 Additional Regularization
While 3DGS optimizes the objective by only sampling the center of each pixel, we found this strategy to be problematic for sparse input setting. In this case, the Gaussians will be deformed to match the color at the center of each pixel, leaving the remaining areas uncovered. As a result, the surfaces, when viewed from a novel view, will appear semi-transparent. To address this issue, we perform the optimization by evaluating multiple samples within each pixel and averaging them to obtain the pixel color. The multisampling ensures that the Gaussians properly cover each pixel, resulting in significantly improved images, as shown in the supplementary video and Table 3.
@@
An alternative approach is to utilize multi-view stereo (MVS) [61, 66, 58] which provides a 3D consistent alternative to monocular depth supervision, as shown by MVSNeRF [10]. However, the state-of-the-art MVS approaches generate depth with significant artifacts for sparse images with large baseline, which hinders the optimization. High quality monocular depth better complements our decoder-based approach to obtain a coherent representation.
@@
where $r$ is the radius of the sphere and is used to create the isometric scale matrix, and $f$ represents the vertical focal length of the input image with height $H$. For the color, represented in terms of the SH coefficients, we set the DC value to the pixel color and other coefficients to zero. The initial opacity of gaussians is set based on the number of views, $2=0.6,3=0.5,4=0.35$ across all datasets. We decrease the initial opacity with the number of views to ensure gradient propagation and better alignment.
@@ 5.1 Comparisons
As discussed previously, unlike other methods that fill in the occluded regions with blurry and repetitive textures, our method does not reconstruct these areas. This gives us a unique advantage as we can identify and inpaint these areas to hallucinate high-quality details. As a proof of concept, we use a simple strategy to inpaint these areas with a diffusion model and project them to 3D using monocular depth (see supplementary Sec.2 for more details). As shown in Fig. 1 (right) the hallucinated details are of high-quality and consistent (see supplementary video).
@@ 6 Conclusion
Acknowledgements
The project was funded in part by a generous gift from Meta.
@@ 7 Implicit Decoder Architecture
We use the convolutional decoder network as proposed by Bemana et al. [5, 37] to output a full resolution 2D map corresponding to the camera index. The normalized scalar input index is concatenated to the coordconv [31] layer in the first layer of the decoder. This is followed by a series of convolutional and upsampling layers (bi-linear interpolation) until the desired output resolution is reached. The number of parameters is controlled using a scalar capacity factor that is multiplied to a preset of each layer, e.g., [2c, 4c, 8c] where c is the capacity factor. We use capacity factor [10, 15, 18] for the depth decoder corresponding to 2, 3 and 4 inputs, respectively. Similarly, we use [6, 10, 12] for the opacity decoder.
8 3D Inpainting
Since we constrain the movement of Gaussians, our approach, unlike existing methods, does not hallucinate details in the occluded areas. This is a unique advantage as it allows us to inpaint these regions using state-of-the-art methods. To do so, we begin by rendering a novel view with holes and alpha mask. We use stable diffusion inpainting [42] to generate the missing texture in the masked regions. Then, we estimate the monocular depth on the inpainted image. Next, we use the gradient of this depth to fill in rendered depth in the masked regions using poisson blending [38]. We then project Gaussians to the scene using the inpainted depth and image. We repeat this process sequentially for a few novel views. Once this process is done, the inpainted scene is 3D consistent and can be rendered from any novel view without holes. We show additional inpainting results in Fig. 11.
9 Comparison against DNGaussian
We compare our approach against a recent state-of-the-art sparse view synthesis approach by Li et al. [27] which also utilizes the 3D Gaussian representation. We use the author provided results for both qualitative (Fig. 12) and quantitative (Table 4) comparisons on the 3-input LLFF [33] dataset. We outperform DNGaussian both visually and numerically since DNGaussian generates blurry texture with distracting artifacts.
10 Additional Ablation Results
To improve the effectiveness of decoder, we estimate a multi-channel depth offset by utilizing a multi-channel depth-based segmentation mask. Table. 5 shows the effectiveness of this multi-channel approach.
Figure 11: 3D consistent scene inpainting results.
Figure 12: LLFF 3 input. We show comparisons against DNGaussian [27].
Table 4: Numerical comparisons with DNGaussian [27] on the LLFF dataset with 3 views.
Method PSNR SSIM LPIPS
DNGaussian 19.55 0.647 0.264
Ours 20.33 0.725 0.180
Table 5: Numerical comparisons to highlight the effect of #channels in the segmentation mask.
Method PSNR SSIM LPIPS
Ours w/o mask 19.70 0.673 0.214
Ours w/ 3ch mask 19.86 0.704 0.198
Ours w/ 5ch mask 20.33 0.725 0.180
Ours w/ 7ch mask 20.13 0.713 0.187